OpenStudy (pgreene008):

Find the cube roots of 8(cos 216° + i sin 216°)

OpenStudy (pgreene008):

fan and medal

OpenStudy (pgreene008):

please help

hartnn (hartnn):

what have you tried?

Directrix (directrix):

I think you'll need this: De Moivre’s Theorem A formula useful for finding powers and roots of complex numbers. @pgreene008

OpenStudy (unklerhaukus):

\[\cos\theta+i\sin\theta=e^{i\theta}\]